In a cylinder there are 60gm of neon and 124gm of oxygen.If pressure of mixture of gases is 30bar.then partial pressure of oxygen in bar?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Pt = Pa + Pb
Pa = Xa*Pt
Pb = (1 – Xa)*Pt
Now moles of the gaes are:
60/20 => 3 moles
and
124/32 =  4 moles 
respectively.
Mole fraction of Ne gas = > 3/(3 + 4) = 3/7 = Xa
Putting in Pa and Pb we get,
Pa = (3/7) * 30 = 90/7 = > 12.84 bar.
Pb = (4/7)*30 = > 120/7 = > 17.14 bar.